What is the change of momentum of a ball weighing a mass 0.2kg rolls with a velocity of 2.0 m/s and rebounds back at 2.0m/s
1
Expert's answer
2020-11-25T04:50:00-0500

We can find the change of momentum of a ball from the definition of the momentum:


"\Delta p=m\Delta v=m(v_f-v_i),""\Delta p=0.2\ kg\cdot(-2.0\ \dfrac{m}{s}-2.0\ \dfrac{m}{s})=-0.8\ \dfrac{kgm}{s}."

The sign minus means that the change of momentum of the ball is directed away from the wall.

Answer:

"\Delta p=0.8\ \dfrac{kgm}{s}," away from the wall.